### v4

# Dual-Regularized Riccati Recursions for Interior-Point Optimal Control

Abstract
We derive closed-form extensions of Riccati’s recursions (both sequential [4] and parallel [7]) for solving dual-regularized LQR problems. We show how these methods can be used to solve general constrained, non-convex, discrete-time optimal control problems via a regularized interior point method, while guaranteeing that each primal step is a descent direction of an Augmented Barrier-Lagrangian merit function. We provide MIT-licensed implementations of our methods in C++ and JAX.
I Introduction
The object of optimal control theory is the study of optimization problems where the cost functional being optimized depends on the state ($x$) and control ($u$) time-series of a certain dynamical system, whose dynamics are described by an ordinary differential equation $\dot{x}(t)=f(x(t),u(t),t)$ (in the case of continuous-time problems) or by a difference equation $x_{n+1}=f_{n}(x_{n},u_{n})$ (in the case of discrete-time problems).
Numerical optimal control, both real-time and offline, has found numerous application domains, ranging from trajectory optimization for robotics (e.g. for autonomous cars, unmanned aerial vehicles, legged robots) and airspace (e.g. rockets, satellites) to the control of power systems (e.g. power plants, electrical grid).
Continuous-time optimal control problems, whose optimization variables are functions (thus infinite-dimensional) are typically converted into finite-dimensional optimization problems by either shooting (i.e. replacing $x(t),u(t)$ with sampled values $(x_{i})_{i=0}^{N},(u_{i})_{i=0}^{N-1}$) or collocation (i.e. expressing the $x(t),u(t)$ as finitely-parameterized functions such as splines). This process is called transcription. When the problem is transcribed via a shooting method and an explicit integration scheme is used to discretize the dynamics, the resulting optimization problem is a discrete-time optimal control problem.
Depending on the application domain, different transcription mechanisms may be more suitable. In robotics, where optimal control problems are often solved at a high frequency, schemes resulting in faster optimization times are often favored to those that may yield more precise solutions at the cost of slower optimization times. The opposite may be true of other application domains, such as offline trajectory optimization for satellites or rockets.
Henceforth, we shall only discuss discrete-time optimal control problems. Solution mechanisms for these problems can be roughly divided into two groups: single-shooting and multiple-shooting methods. In the case of single-shooting, only the control variables $u_{i}$ are free optimization variables, and the state variables $x_{i}$ are treated as dependent variables; in particular, the dynamics of the system are satisfied at every iterate. However, this comes at the cost of making it substantially harder to warm-start the optimization process. The most common single-shooting methods are the Differentiable Dynamic Programming (DDP) algorithm [5] and the Stagewise Newton algorithm [3]. In the case of multiple shooting methods, some or all of the state variables $x_{n+1}$ are also treated as independent variables, and the dynamics equations $x_{n+1}=f_{n}(x_{u},u_{n})$ for such $n$ are enforced as general constraints. This formulation has the advantage of being easy to warm-start, as well as of being easily paired with generic optimizers of nonlinear programs. However, using optimal-control-specific subproblem solvers is crucial for performance.
When the dynamics of the system are affine, the costs are quadratic, and no further constraints exist, the problem can be solved exactly in a single backward and forward pass, as first shown in [4]. The backward pass computes an optimal affine control policy $u_{i}=K_{i}x_{i}+k_{i}$ in $O(N)$ time, and the forward pass alternates between computing the next state via applying the dynamics law and computing the next control by applying the control law. The closed-form equations that produce this backward pass are called the Riccati recursion. The Riccati recursion was first used for solving the subproblems posed by an interior point method in [6].
II Background
II-A Optimal Control
Definition 1
A discrete-time optimal control problem is an optimization problem of the form
$\min_{x_{0},u_{0},\ldots,x_{N}}$ $\sum\limits_{i=0}^{N-1}f_{i}(x_{i},u_{i})+f_{N}(x_{N})$
s.t. $x_{0}=s_{0},$
$\forall i\in\{0,\ldots,N-1\},x_{i+1}=d_{i}(x_{i},u_{i}),$
$\forall i\in\{0,\ldots,N-1\},c_{i}(x_{i},u_{i})=0,$
$\forall i\in\{0,\ldots,N-1\},g_{i}(x_{i},u_{i})\leq 0,$
$c_{N}(x_{N})=0,$
$g_{N}(x_{N})\leq 0.$
The variables $x_{i},u_{i}$ represent the states and controls; the functions $d_{i},f_{i},c_{i},g_{i}$ are the dynamics, costs, equality constraints, and inequality constraints (respectively); $s_{0}$ is the fixed initial state; $N$ is the number of stages.
II-B The Regularized Interior Point Method
The regularized interior point method (IPM) is a way of solving optimization problems of the form
$\min\limits_{x}f(x)\qquad\mbox{s.t.}\quad c(x)=0\wedge g(x)+s=0\wedge s\geq 0,$
where the functions $f,c,g$ are required to be continuously differentiable.
Definition 2
The Barrier-Lagrangian is defined as
$\mathcal{L}(x,s,y,z;\mu)=$ $f(x)-\mu\sum\limits_{i}\log(s_{i})$
$+y^{T}c(x)+z^{T}(g(x)+s)$
Definition 3
The Augmented Barrier-Lagrangian is defined as
$\mathcal{A}(x,s,y,z;\mu,\eta)$
$=$ $\mathcal{L}(x,s,y,z;\mu)+\frac{\eta}{2}(\lVert c(x)\rVert^{2}+\lVert g(x)+s\rVert^{2}).$
At each iteration of the regularized interior point method, the search direction $(\Delta x,\Delta s,\Delta y,\Delta z)$ is computed by solving the linear system
$\begin{bmatrix}P&0&C^{T}&G^{T}\\ 0&S^{-1}Z&0&I\\ C&0&-\frac{1}{\eta}I&0\\ G&I&0&-\frac{1}{\eta}I\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y\\ \Delta z\end{bmatrix}$$\nabla\mathcal{L}$, (1)
where $\nabla\mathcal{L}$ is evaluated at $(x,s,y,z;\mu)$, $P$ is a positive-definite approximation of $\nabla^{2}_{xx}\mathcal{L}(x,s,y,z;\mu)$, $C=J(c)(x)$, $G=J(g)(x)$, and $S,Z$ denote the diagonal matrices with entries $s,z$ respectively.
We call $x,s$ the primal variables and $y,z$ the dual variables. As usual, we require that $s,z>0$ always holds. Below, we will show that primal search direction determined by the regularized interior point method is guaranteed to be a descent direction of $\mathcal{A}(\cdot,\cdot,y,z;\mu,\eta)$ at $(x,s)$ unless the primal variables have already converged.
Lemma 1
The linear system
$\begin{bmatrix}P&0&C^{T}&G^{T}\\ 0&S^{-1}Z&0&I\\ C&0&-\frac{1}{\eta}I&0\\ G&I&0&-\frac{1}{\eta}I\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y-\eta c(x)\\ \Delta z-\eta(g(x)+s)\end{bmatrix}$ (2)
$=$ $-\begin{bmatrix}\nabla_{x}\mathcal{A}(x,s,y,z;\mu,\eta)\\ \nabla_{s}\mathcal{A}(x,s,y,z;\mu,\eta)\\ 0\\ 0\end{bmatrix}$
is equivalent to eq. 1.
Below, we use $D(\cdot;\cdot)$ to represent the directional derivative operator.
Theorem 1
When $\Delta x\neq 0$ or $\Delta s\neq 0$, $D(\mathcal{A}(\cdot,\cdot,y,z;\mu,\eta);(\Delta x,\Delta s))(x,s)<0$.
Proof:
$D(\mathcal{A}(\cdot,\cdot,y,z;\mu,\eta);(\Delta x,\Delta s))(x,s)$
$=$ $\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}\nabla_{x}\mathcal{A}(x,s,y,z;\mu,\eta)\\ \nabla_{s}\mathcal{A}(x,s,y,z;\mu,\eta)\\ \end{bmatrix}$
$=$ $\begin{bmatrix}\Delta x^{T}&\Delta s^{T}&0&0\end{bmatrix}\begin{bmatrix}\nabla_{x}\mathcal{A}(x,s,y,z;\mu,\eta)\\ \nabla_{s}\mathcal{A}(x,s,y,z;\mu,\eta)\\ 0\\ 0\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}&0&0\end{bmatrix}\begin{bmatrix}P&0&C^{T}&G^{T}\\ 0&S^{-1}Z&0&I\\ C&0&-\frac{1}{\eta}I&0\\ G&I&0&-\frac{1}{\eta}I\end{bmatrix}$
$\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y-\eta c(x)\\ \Delta z-\eta(g(x)+s)\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}P&0\\ 0&S^{-1}Z\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\end{bmatrix}$
$-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}C^{T}&G^{T}\\ 0&I\end{bmatrix}\begin{bmatrix}\Delta y-\eta c(x)\\ \Delta z-\eta(g(x)+s)\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}P&0\\ 0&S^{-1}Z\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\end{bmatrix}$
$-\begin{bmatrix}C\Delta x\\ G\Delta x+\Delta s\end{bmatrix}^{T}\begin{bmatrix}\eta(C\Delta x)\\ \eta(G\Delta x+\Delta s)\end{bmatrix}$
$=$ $-\begin{bmatrix}\Delta x^{T}&\Delta s^{T}\end{bmatrix}\begin{bmatrix}P&0\\ 0&S^{-1}Z\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\end{bmatrix}$
$-\eta(\lVert C\Delta x\rVert^{2}+\lVert G\Delta x+\Delta s\rVert^{2})<0$
Thus, the Augmented Barrier-Lagrangian $\mathcal{A}(\cdot,\cdot,y,z;\mu,\eta)$ can be used as the merit function for a line search over the primal variables $(x,s)$.
Finally, note that the variable $\Delta s$ can be eliminated from eq. 1 via
$\Delta s$ $=-Z^{-1}S(\Delta z-\mu S^{-1}e+z)$
$=-Z^{-1}S\Delta z+\mu Z^{-1}e-s,$
where $e$ represents the all-$1$ vector, resulting in the linear system
$\begin{bmatrix}P&C^{T}&G^{T}\\ C&-\frac{1}{\eta}I&0\\ G&0&-(W+\frac{1}{\eta}I)\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta y\\ \Delta z\end{bmatrix}$ (3)
$=$ $-\begin{bmatrix}\nabla_{x}\mathcal{L}(x,s,y,z;\mu,\eta)\\ c(x)\\ g(x)+\mu Z^{-1}e\end{bmatrix}$,
where $W=Z^{-1}S$.
II-C Applying the regularized IPM to optimal control
As shown in section II-B, the line search directions used in the regularized interior point method are computed by solving a linear system of the form
$\begin{bmatrix}P&C^{T}&G^{T}\\ C&-\frac{1}{\eta}I&0\\ G&0&-W-\frac{1}{\eta}I\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta y\\ \Delta z\end{bmatrix}=-\begin{bmatrix}r_{x}\\ r_{y}\\ r_{z}\end{bmatrix}$,
where $\eta>0$ and $W$ is a diagonal matrix with positive elements.
Note that $\Delta z$ can be eliminated via $\Delta z=(W+\frac{1}{\eta}I)^{-1}(G\Delta x+r_{z})$, resulting in
$\begin{bmatrix}P+G^{T}(W+\frac{1}{\eta}I)^{-1}G&C^{T}\\ C&-\frac{1}{\eta}I\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta y\end{bmatrix}$
$=$ $-\begin{bmatrix}r_{x}+G^{T}(W+\frac{1}{\eta}I)^{-1}r_{z}\\ r_{y}\end{bmatrix}$.
Any component of $\Delta y$ corresponding to an equality constraint other than the dynamics can be eliminated in the same fashion. Importantly, note that the only constraints that have cross-stage dependencies are the dynamics, so these eliminations preserve the stagewise nature of the problem.
This leaves us with a linear system that is a Dual-Regularized LQR problem.
II-D Dual-Regularized LQR
Definition 4
A dual-regularized LQR problem is a linear system of the form
$\begin{bmatrix}P&C^{T}\\ C&-\Delta\end{bmatrix}\begin{bmatrix}x\\ y\end{bmatrix}=-\begin{bmatrix}s\\ c\end{bmatrix}$, (4)
where
$P$ $=\begin{bmatrix}P_{0}&&\\ &\ddots&\\ &&P_{N}\end{bmatrix},\quad\Delta=\begin{bmatrix}\delta_{0}&&\\ &\ddots&\\ &&\delta_{N}\end{bmatrix}$, (5)
[…]
$C$ $=\begin{bmatrix}-I&&&&&&\\ A_{0}&B_{0}&-I&&&&\\ &&&\ddots&-I&&\\ &&&&A_{N-1}&B_{N-1}&-I\end{bmatrix}$,
$s$ $=\begin{bmatrix}q_{0}\\ r_{0}\\ \vdots\\ q_{N-1}\\ r_{N-1}\\ q_{N}\end{bmatrix},c=\begin{bmatrix}c_{0}\\ c_{1}\\ \vdots\\ c_{N}\end{bmatrix},x=\begin{bmatrix}x_{0}\\ u_{0}\\ \vdots\\ x_{N-1}\\ u_{N-1}\\ x_{N}\end{bmatrix},y=\begin{bmatrix}y_{0}\\ y_{1}\\ \vdots\\ y_{N}\end{bmatrix}$.
Moreover, the matrices $P_{i}$ are expected to be symmetric and positive semi-definite; the matrices $R_{i}$ are expected to be symmetric and positive definite; the scalars $\delta_{0},\ldots,\delta_{N}$ are expected to be positive.
Note that, when $\Delta=0$, this linear system is the KKT system associated with a standard LQR problem (i.e. not dual-regularized, and viewed as an optimization problem).
Definition 5
The dual-regularized Lagrangian associated with eq. 4 is defined as
$\mathcal{R}(x,y)=\frac{1}{2}x^{T}Px+s^{T}x+y^{T}(Cx+c)-\frac{1}{2}y^{T}\Delta y.$
The linear system from eq. 4 can be seen as the KKT system associated with the optimization problem
$\max\limits_{y}\min\limits_{x}\mathcal{R}(x,y).$ (6)
Specializing eq. 6 to the case of discrete-time optimal control problems, as in eq. 5, it becomes
$\max\limits_{y_{0},\ldots,y_{N}}\min\limits_{x_{0},u_{0},\ldots,x_{N}}\mathcal{R}(x_{0},u_{0},\ldots,x_{N},y_{0},\ldots,y_{N}),$ (7)
where
$\mathcal{R}(x_{0},u_{0},\ldots,x_{N},y_{0},\ldots,y_{N})=$
$\sum\limits_{i=0}^{N-1}\left(\frac{1}{2}x_{i}^{T}Q_{i}x_{i}+x_{i}^{T}M_{i}u_{i}+\frac{1}{2}u_{i}^{T}R_{i}u_{i}+q_{i}^{T}x_{i}+r_{i}^{T}u_{i}\right)$
$+\frac{1}{2}x_{N}^{T}Q_{N}x_{N}+q_{N}^{T}x_{N}+y_{0}^{T}(c_{0}-x_{0})-\sum\limits_{i=0}^{N}\frac{\delta_{i}}{2}y_{i}^{T}y_{i}$
$+\sum\limits_{i=0}^{N-1}\left(y_{i+1}^{T}(A_{i}x_{i}+B_{i}u_{i}+c_{i+1}-x_{i+1})\right).$
By abuse of notation, we also call eq. 7 a dual-regularized LQR problem.
II-E Associative Scans Overview
Associative scans are a common parallelization mechanism used in functional programming, first introduced in [1]. They were used in [7] to derive a simple method for solving (primal) LQR problems in $O(\log^{2}(m)+\log(N)\log^{2}(n))$ parallel time, where $N,n,m$ are respectively the number of stages, states, and controls. Note, however, that [7] states that the complexity of their algorithm is $O(m+\log(N)n)$, apparently stemming from some confusion around the parallel time complexity of matrix inversion (which is $O(\log(n)^{2})$ for $n\times n$ matrices, as shown in [2]).
Given a set $\mathcal{X}$, a function $f:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{X}$ is said to be associative if $\forall a,b,c\in\mathcal{X},f(f(a,b),c)=f(a,f(b,c))$. The forward associative scan operation $S_{f}(x_{1},\ldots,x_{n};f)$ can be inductively defined by
$S_{f}(x_{1};f)$ $=x_{1},$
$S_{f}(x_{1},\ldots,x_{i+1};f)$ $=(y_{1},\ldots,y_{i},f(y_{i},x_{i+1})),$
where $y_{1},\ldots,y_{i}=S_{f}(x_{1},\ldots,x_{i};f)$. Similarly, the reverse associative scan operation $S_{r}(x_{1},\ldots,x_{n};f)$ can be inductively defined by
$S_{r}(x_{1};f)$ $=x_{1},$
$S_{r}(x_{1},\ldots,x_{i+1};f)$ $=(f(x_{1},y_{2}),y_{2},\ldots,y_{i+1}),$
where $y_{2},\ldots,y_{i+1}=S_{r}(x_{2},\ldots,x_{i+1};f)$. [1] provides a method for performing associative scans of $N$ elements in parallel time $O(\log(N))$.
We will employ associative scans do derive the parallel dual-regularized Riccati recursion, effectively extending [7].
III Motivation
The reader may ask why we wish to regularize the dynamics and initial state constraints, whose Jacobians always have full row rank due to the presence of the identity blocks. In fact, if we regularized all equality and inequality constraints except for the dynamics, and performed the variable eliminations described in section II-C, we would be left with a standard LQR problem. The issue with doing so is that the computed primal search direction would no longer be guaranteed to be a descent direction of the Augmented Barrier-Lagrangian with the pre-specified penalty parameter.
IV Sequential Algorithm
We start with some simple (but helpful) lemmas that we will employ to eliminate some of the variables in eq. 7.
Lemma 2
If $\delta>0$ and $f(y)=k^{T}y-\frac{\delta}{2}y^{T}y$, then $\max\limits_{y}f(y)=f(\frac{k}{\delta})=\frac{1}{2\delta}\lVert k\rVert^{2}$.
Proof:
$\nabla f(y)=k-\delta y=0\Rightarrow y=\frac{k}{\delta}$. Moreover, $f(\frac{k}{\delta})=k^{T}\frac{k}{\delta}-\frac{\delta}{2}\frac{k^{T}k}{\delta^{2}}=\frac{k^{T}k}{\delta}-\frac{k^{T}k}{2\delta}=\frac{1}{2\delta}\lVert k\rVert^{2}$. ∎
Lemma 3
If $P$ is symmetric and positive semi-definite and $\delta\geq 0$,
$I-(I+\delta P)^{-1}=(I+\delta P)^{-1}\delta P=\delta P(I+\delta P)^{-1}.$
Proof:
$(I+\delta P)^{-1}\delta P$ $=(I+\delta P)^{-1}(I+\delta P)-(I+\delta P)^{-1}$
$=I-(I+\delta P)^{-1}.$
Similarly,
$\delta P(I+\delta P)^{-1}$ $=(I+\delta P)(I+\delta P)^{-1}-(I+\delta P)^{-1}$
$=I-(I+\delta P)^{-1}.$
Lemma 4
If $P$ is positive semi-definite, $\delta>0$, and $f(x)=\frac{1}{2}x^{T}Px+p^{T}x+\frac{1}{2\delta}\lVert c-x\rVert^{2}$, then
$\min\limits_{x}f(x)=$ $f\left((I+\delta P)^{-1}(c-\delta p)\right)$
$=$ $\frac{1}{2}c^{T}\left((I+\delta P)^{-1}P\right)c+p^{T}c$
$-\delta p^{T}(I+\delta P)^{-1}Pc-\frac{\delta}{2}p^{T}(I+\delta P)^{-1}p.$
Proof:
$\nabla f(x)=Px+p-\frac{c-x}{\delta}=0$
$\Rightarrow$ $(I+\delta P)x=c-\delta p$
$\Rightarrow$ $x=(I+\delta P)^{-1}(c-\delta p).$
Since we can also write
$f(x)=\frac{1}{\delta}\left(\frac{1}{2}x^{T}(I+\delta P)x+(\delta p-c)^{T}x+\frac{1}{2}c^{T}c\right),$
and noting that we use lemma 3 below, it follows that
$f\left((I+\delta P)^{-1}(c-\delta p)\right)$
$=$ $\frac{1}{\delta}\left(\frac{1}{2}(c-\delta p)^{T}(I+\delta P)^{-1}(c-\delta p)\right$.
[…]
$=$ $\frac{1}{2\delta}\left(c^{T}c-(c-\delta p)(I+\delta P)^{-1}(c-\delta p)\right)$
$=$ $\frac{1}{\delta}\left(\frac{1}{2}c^{T}\left(I-(I+\delta P)^{-1}\right)c+\delta p^{T}(I+\delta P)^{-1}c\right$.
$\left.-\frac{\delta^{2}}{2}p^{T}(I+\delta P)^{-1}p\right)$
$=$ $\frac{1}{\delta}\left(\frac{1}{2}c^{T}\left((I+\delta P)^{-1}\delta P\right)c\right$.
$\left.+\delta p^{T}(I-\delta(I+\delta P)^{-1}P)c-\frac{\delta^{2}}{2}p^{T}(I+\delta P)^{-1}p\right)$
$=$ $\frac{1}{2}c^{T}\left((I+\delta P)^{-1}P\right)c+p^{T}c$
$-\delta p^{T}(I+\delta P)^{-1}Pc-\frac{\delta}{2}p^{T}(I+\delta P)^{-1}p.$
For $0\in\{0,\ldots,N\}$, we define the cost-to-go functions
$V_{k}(x_{k})=$ $\max\limits_{y_{k+1},\ldots,y_{N}}\min\limits_{u_{k},x_{k+1},\ldots,u_{N-1},x_{N}}$ (8)
$\sum\limits_{i=k}^{N-1}\left(\frac{1}{2}x_{i}^{T}Q_{i}x_{i}+x_{i}^{T}M_{i}u_{i}+\frac{1}{2}u_{i}^{T}R_{i}u_{i}\right$.
[…]
$+\sum\limits_{i=k}^{N-1}\left(y_{i+1}^{T}\left(A_{i}x_{i}+B_{i}u_{i}+c_{i+1}-x_{i+1}\right)\right)$
$-\sum\limits_{i=k}^{N-1}\frac{\delta_{i+1}}{2}y_{i+1}^{T}y_{i+1}.$
This brings us to the key theorem.
Theorem 2
For all $k\in\{0,\ldots,N\}$, there exist $P_{k},p_{k}$ such that
$\min\limits_{x_{k}}V_{k}(x_{k})=\min\limits_{x_{k}}\left(\frac{1}{2}x_{k}^{T}P_{k}x_{k}+p_{k}^{T}x_{k}\right)+V_{k}(0),$
where $P_{k}$ is symmetric and positive semi-definite.
Moreover, there exist $K_{k},k_{k}$ such that eq. 7 is optimal at
$u_{k}=$ $K_{k}x_{k}+k_{k}$
$x_{0}=$ $(I+\delta_{0}P_{0})^{-1}(c_{0}-\delta_{0}p_{0})$
$x_{k+1}=$ $(I+\delta_{k+1}P_{k+1})^{-1}$
$(A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-\delta_{k+1}p_{k+1})$
$y_{k}=$ $P_{k}x_{k}+p_{k}.$
Proof:
We proceed by induction, in decreasing order of $k$. The base case, consisting of $k=N$, holds trivially by setting $P_{N}=Q_{N}$ and $p_{N}=q_{N}$.
Assuming, by the induction hypothesis, that the statement holds true for $k+1$, we will show it remains true for $k$.
Observing that
$V_{k}(x_{k})=$ $\max\limits_{y_{k+1}}\min\limits_{u_{k},x_{k+1}}V_{k+1}(x_{k+1})+\frac{1}{2}x_{k}^{T}Q_{k}x_{k}$ (9)
$+x_{k}^{T}M_{k}u_{k}+\frac{1}{2}u_{k}^{T}R_{k}u_{k}-\frac{\delta_{k+1}}{2}y_{k+1}^{T}y_{k+1}$
$+y_{k+1}^{T}(A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}),$
we start by eliminating $y_{k+1}$ from eq. 9, by applying lemma 2.
Note that
$\max\limits_{y_{k+1}}y_{k+1}^{T}\left(A_{i}x_{i}+B_{i}u_{i}+c_{k+1}-x_{k+1}\right)-\frac{\delta_{k+1}}{2}y_{k+1}^{T}y_{k+1}$
$=\frac{1}{2\delta_{k+1}}\lVert A_{i}x_{i}+B_{i}u_{i}+c_{k+1}-x_{k+1}\rVert^{2},$
achieved at
$y_{k+1}=\frac{1}{\delta_{k+1}}(A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}).$ (10)
Thus,
$V_{k}(x_{k})=$ $\min\limits_{u_{k},x_{k+1}}V_{k+1}(x_{k+1})+\frac{1}{2}x_{k}^{T}Q_{k}x_{k}+x_{k}^{T}M_{k}u_{k}$
$+\frac{1}{2\delta_{k+1}}\lVert A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}\rVert^{2}$
$+\frac{1}{2}u_{k}^{T}R_{k}u_{k}+\text{const}.$
Applying the induction hypothesis,
$V_{k}(x_{k})=\min\limits_{u_{k},x_{k+1}}\frac{1}{2}x_{k}^{T}Q_{k}x_{k}+x_{k}^{T}M_{k}u_{k}+\frac{1}{2}u_{k}^{T}R_{k}u_{k}$ (11)
$+\frac{1}{2}x_{k+1}P_{k+1}x_{k+1}+p_{k+1}^{T}x_{k+1}$
$+\frac{1}{2\delta_{k+1}}\lVert A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}\rVert^{2}+\text{const}.$
Next, we apply lemma 4 to eliminate $x_{k+1}$ from LABEL:before-eliminating-x.
The terms involving $x_{k+1}$ are
$\min\limits_{x_{k+1}}$ $\frac{1}{2}x_{k+1}^{T}P_{k+1}x_{k+1}+p_{k+1}^{T}x_{k+1}$
$+\frac{1}{2\delta_{k+1}}\lVert A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1}\rVert^{2}.$
Letting
$W_{k+1}$ $=(I+\delta_{k+1}P_{k+1})^{-1}P_{k+1},$
applying lemma 4, and discarding additive constant terms, this becomes
$\frac{1}{2}(A_{k}x_{k}+B_{k}u_{k}+c_{k+1})^{T}W_{k+1}(A_{k}x_{k}+B_{k}u_{k}+c_{k+1})$
$+p_{k+1}^{T}(A_{k}x_{k}+B_{k}u_{k}+c_{k+1})$
$-\delta_{k+1}p_{k+1}^{T}W_{k+1}(A_{k}x_{k}+B_{k}u_{k}+c_{k+1}),$
achieved at
$x_{k+1}=$ $(I+\delta_{k+1}P_{k+1})^{-1}$ (12)
$(A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-\delta_{k+1}p_{k+1}).$
Next, we eliminate $u_{k}$ from the resulting version of LABEL:before-eliminating-x after $x_{k+1}$ has been eliminated. The terms involving $u_{k}$ are
$\min\limits_{u_{k}}(r_{k}+M_{k}^{T}x_{k})^{T}u_{k}+\frac{1}{2}u_{k}^{T}R_{k}u_{k}$ (13)
$+\frac{1}{2}(A_{k}x_{k}+B_{k}u_{k}+c_{k+1})^{T}W_{k+1}(A_{k}x_{k}+B_{k}u_{k}+c_{k+1})$
$+p_{k+1}^{T}B_{k}u_{k}-\delta_{k+1}p_{k+1}^{T}W_{k+1}B_{k}u_{k}$
$=\min\limits_{u_{k}}\frac{1}{2}u_{k}^{T}(R_{k}+B_{k}^{T}W_{k+1}B_{k})u_{k}$
$+(r_{k}+M_{k}^{T}x_{k}+B_{k}^{T}W_{k+1}(A_{k}x_{k}+c_{k+1})$
$+B_{k}^{T}(p_{k+1}-\delta_{k+1}W_{k+1}p_{k+1})^{T}u_{k}$
$+\frac{1}{2}(A_{k}x_{k}+c_{k+1})^{T}W_{k+1}(A_{k}x_{k}+c_{k+1}).$
Note that this leaves out the terms
$p_{k+1}^{T}(A_{k}x_{k}+c_{k+1})-\delta_{k+1}p_{k+1}(W_{k+1}A_{k}x_{k}+c_{k+1}),$ (14)
which do not depent on $u_{k}$. We will pick these back up later.
Letting
$G_{k}$ $=R_{k}+B_{k}^{T}W_{k+1}B_{k}$
$g_{k}$ $=p_{k}+W_{k}(c_{k}-\delta_{k}p_{k})$
$H_{k}$ $=B_{k}^{T}W_{k+1}A_{k}+M_{k}^{T}$
$h_{k}$ $=r_{k}+B_{k}^{T}g_{k+1}$
$K_{k}$ $=-G_{k}^{-1}H_{k}$
$k_{k}$ $=-G_{k}^{-1}h_{k},$
LABEL:only-u-terms becomes
$\min\limits_{u_{k}}$ $\frac{1}{2}u_{k}^{T}G_{k}u_{k}+(H_{k}x_{k}+h_{k})^{T}u_{k}$
$+\frac{1}{2}(A_{k}x_{k}+c_{k+1})^{T}W_{k+1}(A_{k}x_{k}+c_{k+1}).$
Taking the gradient with respect to $u_{k}$ and equating it to 0, we get
$u_{k}=-G_{k}^{-1}(H_{k}x_{k}+h_{k})=K_{k}x_{k}+k_{k}.$ (15)
Plugging this back in, and discarding additive constants, we get
$-\frac{1}{2}(H_{k}x_{k}+h_{k})^{T}G_{k}^{-1}(H_{k}x_{k}+h_{k})$ (16)
$+\frac{1}{2}(A_{k}x_{k}+c_{k+1})^{T}W_{k+1}(A_{k}x_{k}+c_{k+1})$
$=$ $\frac{1}{2}x_{k}(A_{k}^{T}W_{k+1}A_{k}-H_{k}G_{k}^{-1}H_{k})x_{k}$
$+(A_{k}^{T}W_{k+1}c_{k+1}-H_{k}^{T}G_{k}^{-1}h_{k})x_{k}$
$=$ $\frac{1}{2}x_{k}(A_{k}^{T}W_{k+1}A_{k}+H_{k}K_{k})x_{k}$
$+(A_{k}^{T}W_{k+1}c_{k+1}+H_{k}^{T}k_{k})x_{k}.$
Collecting all remaining terms from eq. 14 and LABEL:after-eliminating-u, discarding additive constants, and letting
$P_{k}$ $=A_{k}^{T}W_{k+1}A_{k}+Q_{k}+H_{k}^{T}K_{k}$ (17)
$p_{k}$ $=q_{k}+A_{k}^{T}g_{k+1}+H_{k}^{T}k_{k},$
LABEL:before-eliminating-x becomes
$V_{k}(x_{k})$ $=\frac{1}{2}x_{k}^{T}P_{k}x_{k}+p_{k}^{T}x_{k}+\text{const}$
$=\frac{1}{2}x_{k}^{T}P_{k}x_{k}+p_{k}^{T}x_{k}+V_{k}(0).$
Next, we must show that $P_{k}$ remains positive semi-definite.
Noting that $P_{k}$ is the Schur complement of
$\begin{bmatrix}Q_{k}+A_{k}^{T}W_{k+1}A_{k}&H_{k}^{T}\\ H_{k}&G_{k}\end{bmatrix}$ (18)
$=$ $\begin{bmatrix}Q_{k}&M_{k}^{T}\\ M_{k}^{T}&R_{k}\end{bmatrix}+\begin{bmatrix}A_{k}^{T}W_{k+1}A_{k}&A_{k}^{T}W_{k+1}B_{k}\\ B_{k}^{T}W_{k+1}A_{k}&B_{k}^{T}W_{k+1}B_{k}\end{bmatrix}$
$=$ $\begin{bmatrix}Q_{k}&M_{k}^{T}\\ M_{k}^{T}&R_{k}\end{bmatrix}+\begin{bmatrix}A_{k}^{T}\\ B_{k}^{T}\end{bmatrix}W_{k+1}\begin{bmatrix}A_{k}&B_{k}\end{bmatrix}$,
it follows that $P_{k}$ is positive semi-definite, as $G_{k}$ is positive-definite and the block-matrix LABEL:psd-block-matrices is positive semi-definite.
Noting that eq. 7 can be written as
$\max\limits_{y_{0}}\min\limits_{x_{0}}V_{0}(x_{0})+y_{0}^{T}(x_{0}-c_{0})-\frac{\delta_{0}}{2}y_{0}^{T}y_{0},$
we can apply lemma 2 to re-write this as
$\min\limits_{x_{0}}V_{0}(x_{0})+\frac{1}{2\delta_{0}}\lVert c_{0}-x_{0}\rVert^{2}$
$=\frac{1}{\delta_{0}}\left(\min\limits_{x_{0}}\frac{1}{2}x_{0}^{T}(I+\delta_{0}P_{0})x_{0}+(\delta_{0}p_{0}-c_{0})^{T}x_{0}+c_{0}^{T}c_{0}\right)$
$+V_{0}(0),$
achieved at
$y_{0}=\frac{1}{\delta_{0}}(c_{0}-x_{0}).$ (19)
Solving for $x_{0}$, we get
$x_{0}=(I+\delta_{0}P_{0})^{-1}(c_{0}-\delta_{0}p_{0}).$ (20)
The remaining $x_{k},u_{k}$ can be recovered in a forward pass, i.e. in increasing order of $k$ via eq. 12 and eq. 15.
Finally, we can re-write eq. 19 and eq. 10 to remove any division by $\delta_{0},\ldots,\delta_{N}$, in the interest of improving numerical stability when $\delta_{0},\ldots,\delta_{N}\rightarrow 0$ as well as extending this method to the case of $\delta_{0},\ldots,\delta_{N}=0$ (recovering the standard backward and forward passses of the standard LQR algorithm).
Since, due to eq. 20,
$(I+\delta_{0}P_{0})x_{0}=-(\delta_{0}p_{0}-c_{0})\Leftrightarrow\frac{1}{\delta_{0}}(c_{0}-x_{0})=P_{0}x_{0}+p_{0},$
we can re-write eq. 19 as
$y_{0}=P_{0}x_{0}+p_{0}.$
Moreover, due to eq. 12,
$(I+\delta_{k+1}P_{k+1})x_{k+1}=A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-\delta_{k+1}p_{k+1}$
$\Leftrightarrow$ $\frac{1}{\delta_{k+1}}(A_{k}x_{k}+B_{k}u_{k}+c_{k+1}-x_{k+1})=P_{k+1}x_{k+1}+p_{k+1}.$
Thus, we can re-write eq. 10 as
$y_{k+1}=P_{k+1}x_{k+1}+p_{k+1}.$
V Parallel Algorithm
First, we can eliminate the control variables $u_{i}$ from eq. 7. Computing the gradients with respect to $u_{i}$, equating them to 0, and solving for $u_{i}$ yields
$u_{i}=-R_{i}^{-1}\left(M_{i}^{T}x_{i}+r_{i}+B_{i}^{T}y_{i+1}\right).$
Noting that
$\frac{1}{2}u_{i}^{T}R_{i}u_{i}=$ $\frac{1}{2}x_{i}^{T}M_{i}R_{i}^{-1}M_{i}^{T}x_{i}+r_{i}^{T}R_{i}^{-1}M_{i}^{T}x_{i}$
$+\frac{1}{2}y_{i+1}^{T}B_{i}R_{i}^{-1}B_{i}^{T}y_{i+1}+y_{i+1}^{T}B_{i}R_{i}^{-1}M_{i}^{T}x_{i}$
$+r_{i}^{T}R_{i}^{-1}B_{i}^{T}y_{i+1}+\frac{1}{2}r_{i}^{T}R_{i}^{-1}r_{i},$
$x_{i}^{T}M_{i}u_{i}=$ $-x_{i}^{T}M_{i}R_{i}^{-1}M_{i}^{T}x_{i}-r_{i}^{T}R_{i}^{-1}M_{i}^{T}x_{i}$
$-y_{i+1}^{T}B_{i}R_{i}^{-1}M_{i}^{T}x_{i},$
$r_{i}^{T}u_{i}=$ $-r_{i}^{T}R_{i}^{-1}M_{i}^{T}x_{i}-r_{i}^{T}R_{i}^{-1}r_{i}$
$-y_{i+1}^{T}B_{i}R_{i}^{-1}r_{i},$
$y_{i+1}^{T}B_{i}u_{i}=$ $-y_{i+1}^{T}B_{i}R_{i}^{-1}B_{i}^{T}y_{i+1}$
$-y_{i+1}^{T}B_{i}R_{i}^{-1}M_{i}^{T}x_{i}-r_{i}^{T}R_{i}^{-1}B_{i}^{T}y_{i+1},$
and discarding constant terms, the optimization problem eq. 7 becomes
$\max\limits_{y_{0},\ldots,y_{N}}\min\limits_{x_{0},\ldots,x_{N}}\sum\limits_{i=0}^{N-1}\left(\frac{1}{2}x_{i}^{T}(Q_{i}-M_{i}R_{i}^{-1}M_{i}^{T})x_{i}\right$. (21)
[…]
$\left.\left.+\left(c_{i+1}-B_{i}R_{i}^{-1}r_{i}\right)-x_{i+1}\right)\right)$
$-\frac{1}{2}\sum\limits_{i=0}^{N-1}y_{i+1}^{T}\left(\delta_{i+1}I+B_{i}R_{i}^{-1}B_{i}^{T}\right)y_{i+1}$
$+y_{0}^{T}(c_{0}-x_{0})-\frac{\delta_{0}}{2}y_{0}^{T}y_{0}.$
For $i\in\{0,\ldots,N-1\}$, we define
$P_{i\rightarrow i+1}$ $=Q_{i}-M_{i}R_{i}^{-1}M_{i}^{T},$ (22)
$p_{i\rightarrow i+1}$ $=q_{i}-M_{i}R_{i}^{-1}r_{i},$
$A_{i\rightarrow i+1}$ $=A_{i}-B_{i}R_{i}^{-1}M_{i}^{T},$
$C_{i\rightarrow i+1}$ $=\delta_{i+1}I+B_{i}R_{i}^{-1}B_{i}^{T},$
$c_{i\rightarrow i+1}$ $=c_{i+1}-B_{i}R_{i}^{-1}r_{i},$
Similarly, we define
$P_{N\rightarrow N+1}$ $=Q_{N},$ (23)
$p_{N\rightarrow N+1}$ $=q_{N},$
$A_{N\rightarrow N+1}$ $=0,$
$C_{N\rightarrow N+1}$ $=0,$
$c_{N\rightarrow N+1}$ $=0,$
For $0\leq i<j\leq N+1$, we define the interval value functions
$V_{i\rightarrow j}(x_{i},x_{j})=$ $\max\limits_{y_{i+1},\ldots,y_{j}}\min\limits_{x_{i+1},\ldots,x_{j-1}}$
$\sum\limits_{k=i}^{j-1}\left(\frac{1}{2}x_{k}^{T}P_{k\rightarrow k+1}x_{k}+p_{k\rightarrow k+1}^{T}x_{k}\right)$
$+$ $\sum\limits_{k=i}^{j-1}\left(y_{k+1}^{T}\left(A_{k\rightarrow k+1}x_{k}+c_{k\rightarrow k+1}-x_{k+1}\right)\right)$
- $\frac{1}{2}\sum\limits_{k=i}^{j-1}y_{k+1}^{T}C_{k\rightarrow k+1}y_{k+1}$
Note that, by definition,
$V_{i}(x_{i})=\min\limits_{x_{N+1}}V_{i\rightarrow N+1}(x_{i},x_{N+1}).$ (24)
Since
$\max\limits_{y_{N+1}}\min\limits_{x_{N+1}}-y_{N+1}^{T}x_{N+1}=0,$
achieved by forcing $x_{N+1}=0$, it follows that
$V_{i}(x_{i})=V_{i\rightarrow N+1}(x_{i},0).$ (25)
It was shown in [7] that the functions $V_{i\rightarrow j}$ admit representations of the form
$V_{i\rightarrow j}(x_{i},x_{j})=\max\limits_{y}\left(\frac{1}{2}x_{i}^{T}P_{i\rightarrow j}x_{i}+p_{i\rightarrow j}^{T}x_{i}\right$.
$\left.-\frac{1}{2}y^{T}C_{i\rightarrow j}y+y^{T}\left(A_{i\rightarrow j}x_{i}+c_{i\rightarrow j}-x_{j}\right)\right),$
modulo constant additive terms that are independent of $x_{i},x_{j}$ and can thus be discarded.
The following combination rules, established in [7], can be applied to compute $V_{i\rightarrow k}$ from $V_{i\rightarrow j}$ and $V_{j\rightarrow k}$:
$P_{i\rightarrow k}=$ $A_{i\rightarrow j}^{T}\left(I+P_{j\rightarrow k}C_{i\rightarrow j}\right)^{-1}P_{j\rightarrow k}A_{i\rightarrow j}+P_{i\rightarrow j},$ (26)
$p_{i\rightarrow k}=$ $A_{i\rightarrow j}^{T}\left(I+P_{j\rightarrow k}C_{i\rightarrow j}\right)^{-1}\left(p_{j\rightarrow k}+P_{j\rightarrow k}c_{i\rightarrow j}\right)$
$+p_{i\rightarrow j},$
$A_{i\rightarrow k}=$ $A_{j\rightarrow k}\left(I+C_{i\rightarrow j}P_{j\rightarrow k}\right)^{-1}A_{i\rightarrow j},$
$C_{i\rightarrow k}=$ $A_{j\rightarrow k}\left(I+C_{i\rightarrow j}P_{j\rightarrow k}\right)^{-1}C_{i\rightarrow j}A_{j\rightarrow k}^{T}+C_{j\rightarrow k},$
$c_{i\rightarrow k}=$ $A_{j\rightarrow k}\left(I+C_{i\rightarrow j}P_{j\rightarrow k}\right)^{-1}\left(c_{i\rightarrow j}-C_{i\rightarrow j}p_{j\rightarrow k}\right)$
$+c_{j\rightarrow k}.$
Note that we depart from [7] in eq. 22, but not in eq. 26 or eq. 23.
A reverse associative scan [1] can be used to compute the $P_{i\rightarrow N+1},p_{i\rightarrow N+1}$ (i.e. the $P_{i},p_{i}$ in eq. 17) in $O(\log(N)\log(n)^{2})$.
Next, the $(I+\delta_{i}P_{i})^{-1}$ can be computed in $O(\log(n)^{2})$ parallel time (i.e. $O(1)$ with respect to $N$).
We can then compute $x_{0}$ in $O(\log(n))$ parallel time via eq. 20 and the $K_{i},k_{i}$ from eq. 15 in $O(\log(n)+\log(m)^{2})$ parallel time (i.e. $O(1)$ with respect to $N$ in both cases).
Finally, we wish to compute the $u_{i},x_{i+1}$ from the $K_{i},k_{i}$ via eq. 12 and eq. 15. Note that the sequential LQR forward pass has $O(N)$ parallel time complexity. However, as done in [7], we can reduce the computation of the $x_{i}$ to a sequential composition of affine functions, which can also be parallelized with an associative scan [1] in $O(\log(N)\log(n))$ parallel time. This will be described in section V-A. Due to theorem 2,
$x_{i+1}=$ $(I+\delta_{i}P_{i+1})^{-1}\left(A_{i}x_{i}+B_{i}u_{i}+c_{i+1}-\delta_{i+1}p_{i+1}\right)$
$=$ $(I+\delta_{i+1}P_{i+1})^{-1}$
$\left(A_{i}x_{i}+B_{i}\left(K_{i}x_{i}+k_{i}\right)+c_{i+1}-\delta_{i+1}p_{i+1}\right)$
$=$ $(I+\delta_{i+1}P_{i+1})^{-1}\left(A_{i}+B_{i}K_{i}\right)x_{i}$
$+(I+\delta_{i+1}P_{i+1})^{-1}(B_{i}k_{i}+c_{i+1}-\delta p_{i+1}).$
Once these affine functions have been composed, they can be independently applied to $x_{0}$ to recover all the $x_{i}$ in $O(\log(n))$ parallel time. The $u_{i}$ can then be computed in $O(\log(m)+\log(n)))$ parallel time by independently evaluating $u_{i}=K_{i}x_{i}+k_{i}$. Note that both of these parallel times are $O(1)$ with respect to $N$.
Thus, the combined parallel time complexity of our method is $O(\log(m)^{2}+\log(N)\log(n)^{2})$.
V-A Composing Affine Functions with Associative Scans
In this section, we describe an algorithm for composing $N$ affine functions $F_{i}(x)=M_{i}x_{i}+m_{i}$, as done in [7], with $O(\log(N)\log(n))$ parallel time.
Letting $\mathcal{X}=\mathbb{R}^{n}\times\mathbb{R}^{n\times n}$ and letting $f:\mathcal{X}\rightarrow\mathcal{X}$ be defined by $f((a,B),(c,D))=(Da+c,DB)$, we claim that $f$ is associative. This is simple to verify:
$f(f((a,B),(c,D)),(e,F))$
$=$ $f((Da+c,DB),(e,F))$
$=$ $(F(Da+c)+e,F(DB))$
$=$ $((FD)a+Fc+e,(FD)B)$
$=$ $f((a,B),(Fc+e,FD))$
$=$ $f((a,B),f((c,D),(e,F))).$
Moreover, note that $f$ is the affine function composition operator, as $D(Bx+a)+c=(DB)x+(Da+c)$. Thus, it suffices to apply a forward associative scan to $f$ to obtain all compositions $F_{0},F_{1}\circ F_{0},\ldots,F_{N-1}\circ\cdots\circ F_{0}$ in $O(\log(N)\log(n))$ parallel time.
VI Residual Computation
The residuals of the Newton-KKT linear systems from eq. 4 (when specialized to discrete-time optimal control problems as in eq. 5) are given by
$\begin{bmatrix}\begin{bmatrix}Q_{i}x_{i}+M_{i}u_{i}+A_{i}^{T}y_{i+1}+q_{i}-y_{i}\\ M_{i}^{T}x_{i}+R_{i}u_{i}+B_{i}^{T}y_{i+1}+r_{i}\\ \end{bmatrix}_{i=0,\ldots,N-1}\\ Q_{N}x_{N}+q_{N}-y_{N}\\ c_{0}-\delta_{0}y_{0}-x_{0}\\ \begin{bmatrix}A_{i}x_{i}+B_{i}u_{i}+c_{i+1}-\delta_{i+1}y_{i+1}-x_{i+1}\end{bmatrix}_{i=0,\ldots,N-1}\end{bmatrix}$
Thus, they can be computed in $O(\log(m)+\log(n))$ parallel time. Residual computation is only necessary when applying iterative refinement to solve  eq. 4 to higher precision.
VII Software Contributions
We provide JAX implementations of both the sequential and parallel dual-regularized LQR algorithms [12], including unit tests verifying the correctness of the solutions provided by our methods on a set of random examples satisfying the required definiteness properties.
Modern interior-point solvers, such as IPOPT [15], while supporting several generic linear system solvers, are not designed to easily accommodate linear solvers that specialize to specific problem types (e.g. optimal control problems). In fact, until recently, IPOPT installations did not include all of the required headers for users to integrate any custom user-side linear system solvers, showing that this was not being done at all.
In order to address this limitation, we released a simple regularized interior point solver in [8], which allows users to easily plug in callbacks specializing certain operations to their specific problem type, including:
  • a KKT system factorization callback;
  • a KKT system solve callback;
  • a KKT system residual computation callback.
This effectively splits the solver into a shared backend and a problem-specific frontend. We do exactly that for the case of optimal control in [13].
Adding to this, we also implemented an integration with QDLDL [14] supporting arbitrary user-provided KKT permutations (for optimal fill-in prevention) in [10], resulting in an efficient sparse interior point solver that avoids dense operations entirely. Alternatively, a sparse linear algebra code generation library, such as SLACG [11], can be used to solve the KKT systems and compute their residuals; this often achieves substantial speed-ups for solving sparse problems, as no index computations are performed at runtime.
Examples can be found in [9]. In all cases, with correct usage, dynamic memory allocation is entirely avoided.
All of the provided packages are free and open-source (MIT-licensed).
VIII Conclusion
We derived extensions of the sequential and parallel Riccati Recursions for dual-regularized LQR problems, allowing us to easily and efficiently numerically solve constrained non-linear discrete-time optimal control problems via the regularized interior point method. We also provided free, efficient, open-source implementations of the described algorithms, which are fully unit-tested on random well-defined examples, establishing the correctness of our derivations and implementations.